# HawkBench: Investigating Resilience of RAG Methods on Stratified Information-Seeking Tasks

### v1

@@ 3.3 Key Insights
Current RAG methods tend to be optimized for specific types of information-seeking tasks (e.g., fact retrieval or rationale generation). However, this specialization leads to a lack of overall resilience across a broader range of tasks. While empirical analyses provide heuristics to guide method selection for particular tasks, we still lack a systematic, adaptable solution that can handle diverse tasks with varying requirements. This gap emphasizes the need for developing RAG systems that can dynamically adjust to different information-seeking challenges, moving beyond task-specific optimizations toward a more generalized framework.
Global Awareness: Construction and Utilization Challenges
@@
Global awareness is essential for tasks that require the integration of information from multiple sources. However, current global RAG methods struggle with efficiently building and fully leveraging this awareness. While methods such as GraphRAG (which uses graph construction) and memory-based approaches show promise, their reliance on inefficient global intermediate construction processes (e.g., building graphs or memory stores) remains a major bottleneck. For example, graph construction can take tens of minutes, making it impractical for real-time use. Optimizing these construction processes could make these systems more viable. Additionally, there is a need for research into how to best utilize global intermediates (e.g., graphs, memory caches) to improve retrieval and reasoning. Exploring efficient ways to construct and use these intermediates is an important direction for future work.
Dynamic Task Understanding and Adaptive Query Interpretation
@@
As information-seeking tasks become more complex, the need for dynamic task understanding and adaptive query interpretation becomes increasingly important. A one-size-fits-all solution is not feasible; instead, RAG systems must integrate decision-making mechanisms that allow them to dynamically adjust how they access (referencing) and utilize (reasoning) knowledge. By understanding the task context and adapting the retrieval strategy accordingly, RAG systems can more effectively address a wider range of queries. This adaptability would significantly enhance the robustness and efficiency of RAG methods, enabling them to handle varying complexities and task types more effectively.
The Potential of Agentic Information-Seeking Systems
Looking ahead, agentic information-seeking systems—capable of autonomously navigating knowledge acquisition—represent a promising frontier. These systems could perform complex tasks, such as writing papers or conducting literature surveys, by integrating retrieval, reasoning, and synthesis. Recent advancements, such as OpenAI’s Deep Research, highlight the potential for these agentic systems to become next-generation solutions for a wide range of tasks. With the ability to autonomously manage complex information-seeking tasks, these systems could transform how we approach knowledge-intensive tasks, making them an exciting area for future exploration.
@@ 4 Related Work
RAG was introduced by Lewis et al. (2020) to enhance language models’ ability to handle knowledge-intensive tasks by providing relevant context through retrieval. Research in RAG has focused on two main areas: (1) improving retrieval quality to set an upper bound for generation accuracy Qian et al. (2024a); Gao et al. (2024), and (2) optimizing the use of retrieved passages for relevance and accessibility during generation Jiang et al. (2023); Zhao et al. (2024a).
The integration of RAG with LLMs has gained momentum, especially in knowledge-intensive applications Shuster et al. (2021). As a result, there is increasing demand for more generalized RAG systems capable of handling a wider range of tasks, including those beyond factoid queries Zhao et al. (2024b). However, traditional RAG pipelines face challenges in addressing complex tasks with implicit information needs, often failing to provide sufficient context for accurate generation Gao et al. (2024); Zhao et al. (2024b). Recent advances have aimed to expand RAG’s applicability. For example, GraphRAG Edge et al. (2024) and HippoRAG Gutiérrez et al. (2024) introduce knowledge graphs to facilitate retrieval and enhance global awareness. Agent-based approaches, such as ActiveRAG Xu et al. (2024); Yoon et al. (2024), plan information access and utilization via agents.
@@
As RAG systems are increasingly adopted, the need for comprehensive evaluation benchmarks has become evident. Early benchmarks, such as KILT Petroni et al. (2021), primarily focused on task-specific aspects like single-hop and multi-hop reasoning, as well as factoid queries. Recently, new benchmarks have been developed to address specialized tasks and domains. For example, MultiHop-RAG evaluates multi-hop tasks Tang and Yang (2024), LegalBench-RAG focuses on the legal domain Pipitone and Alami (2024), CRAG offers a comprehensive evaluation framework for factoid question answering tasks, and RAGBench is designed to assess the explainability of RAG systems Friel et al. (2024). While these benchmarks provide valuable insights into various facets of RAG performance, they lack a comprehensive framework to evaluate the resilience of RAG systems when faced with diverse information-seeking needs, particularly for stratified queries Zhao et al. (2024b).
@@ 5 Conclusion
In this paper, we introduce HawkBench, a comprehensive framework designed to evaluate the resilience of RAG systems across diverse information-seeking tasks. HawkBench is distinguished by its systematic task stratification, multi-domain corpora, and high-quality annotations, making it an robust tool for assessing the resilience of RAG methods. Our evaluation of representative RAG methods reveals that while current RAG systems are often optimized for specific tasks, they lack resilience across general tasks. This highlights the need for dynamic task strategies that integrate decision-making, query interpretation, and global knowledge utilization to enhance the generalizability of RAG systems. HawkBench thus serves as a critical resource for advancing the development of resilient, versatile RAG systems capable of addressing a wide range of real-world user needs.
Limitations
@@ Appendix A Implementation Details
In our evaluation of baseline methods on HawkRAG, we use BGE-M3 Chen et al. (2023) as the retriever for vanilla RAG, RQ-RAG, HyDE, and MemoRAG, setting the hit number to 5. For methods that segment long contexts into chunks, we utilize the semantic-text-splitter tool, limiting chunks to a maximum of 512 tokens. MemoRAG employs the officially released memorag-qwen2-7b-inst as its memory model. For GraphRAG, we leverage GPT-4o for graph construction and use the retrieved context for generation. All baseline methods adopt Qwen-2.5-7B-instruct-128K as the generator.
@@
Table 5: Full details of main experimental results.

### v2

@@ 3.3 Key Insights
Current RAG methods Lack Resilience. Current RAG methods tend to be optimized for specific types of information-seeking tasks (e.g., fact retrieval or rationale generation). However, this specialization leads to a lack of overall resilience across a broader range of tasks. While empirical analyses provide heuristics to guide method selection for particular tasks, we still lack a systematic, adaptable solution that can handle diverse tasks with varying requirements. This gap emphasizes the need for developing RAG systems that can dynamically adjust to different information-seeking challenges, moving beyond task-specific optimizations toward a more generalized framework.
@@
Global Awareness: Construction and Utilization Challenges. Global awareness is essential for tasks that require the integration of information from multiple sources. However, current global RAG methods struggle with efficiently building and fully leveraging this awareness. While methods such as GraphRAG (which uses graph construction) and memory-based approaches show promise, their reliance on inefficient global intermediate construction processes (e.g., building graphs or memory stores) remains a major bottleneck. For example, graph construction can take tens of minutes, making it impractical for real-time use. Optimizing these construction processes could make these systems more viable. Additionally, there is a need for research into how to best utilize global intermediates (e.g., graphs, memory caches) to improve retrieval and reasoning. Exploring efficient ways to construct and use these intermediates is an important direction for future work.
@@
Dynamic Task Understanding and Adaptive Query Interpretation. As information-seeking tasks become more complex, the need for dynamic task understanding and adaptive query interpretation becomes increasingly important. A one-size-fits-all solution is not feasible; instead, RAG systems must integrate decision-making mechanisms that allow them to dynamically adjust how they access (referencing) and utilize (reasoning) knowledge. By understanding the task context and adapting the retrieval strategy accordingly, RAG systems can more effectively address a wider range of queries. This adaptability would significantly enhance the robustness and efficiency of RAG methods, enabling them to handle varying complexities and task types more effectively.
The Potential of Agentic Information-Seeking Systems. Looking ahead, agentic information-seeking systems—designed to autonomously navigate knowledge acquisition—offer a compelling direction for the future of AI. By integrating retrieval, reasoning, and synthesis, these systems can perform complex tasks such as literature reviews, report writing, or exploratory research. Recent developments like OpenAI’s Deep Research exemplify this trend, signaling a shift toward AI agents that not only assist but independently manage knowledge-intensive workflows. As these systems mature, they hold the potential to reshape how we interact with and generate information, making them a key area for future investigation and innovation.
@@ 4 Related Work
RAG was introduced by Lewis et al. (2020) to enhance language models’ ability to handle knowledge-intensive tasks by providing relevant context through retrieval. Research in RAG has focused on two main areas: (1) improving retrieval quality to set an upper bound for generation accuracy (Qian et al., 2024; Gao et al., 2024), and (2) optimizing the use of retrieved passages for relevance and accessibility during generation (Jiang et al., 2023; Zhao et al., 2024b).
The integration of RAG with LLMs has gained momentum, especially in knowledge-intensive applications (Shuster et al., 2021). As a result, there is increasing demand for more generalized RAG systems capable of handling a wider range of tasks, including those beyond factoid queries (Zhao et al., 2024a). However, traditional RAG pipelines face challenges in addressing complex tasks with implicit information needs, often failing to provide sufficient context for accurate generation (Gao et al., 2024; Zhao et al., 2024a). Recent advances have aimed to expand RAG’s applicability. For example, GraphRAG (Edge et al., 2024) and HippoRAG (Jimenez Gutierrez et al., 2024) introduce knowledge graphs to facilitate retrieval and enhance global awareness. Agent-based approaches, such as ActiveRAG (Xu et al., 2024; Yoon et al., 2024), plan information access and utilization via agents.
@@
As RAG systems are increasingly adopted, the need for comprehensive evaluation benchmarks has become evident. Early benchmarks, such as KILT (Petroni et al., 2021), primarily focused on task-specific aspects like single-hop and multi-hop reasoning, as well as factoid queries. Recently, new benchmarks have been developed to address specialized tasks and domains. For example, MultiHop-RAG evaluates multi-hop tasks (Tang and Yang, 2024), LegalBench-RAG focuses on the legal domain (Guha et al., 2023), CRAG offers a comprehensive evaluation framework for factoid question answering tasks, and RAGBench is designed to assess the explainability of RAG systems (Friel et al., 2024). While these benchmarks provide insights into various facets of RAG performance, they lack a comprehensive framework to evaluate the resilience of RAG systems when faced with diverse information-seeking needs, particularly for stratified queries (Zhao et al., 2024a).
@@ 5 Conclusion
In this paper, we introduce HawkBench, a comprehensive framework designed to evaluate the resilience of RAG systems across diverse information-seeking tasks. HawkBench is distinguished by its systematic task stratification, multi-domain corpora, and high-quality annotations, making it an robust tool for assessing the resilience of RAG methods. Our evaluation of representative RAG methods reveals that while current RAG systems are often optimized for specific tasks, they lack resilience across general tasks. This highlights the need for dynamic task strategies that integrate decision-making, query interpretation, and global knowledge utilization to enhance the generalizability of RAG systems. HawkBench serves as a critical resource for advancing the development of resilient, versatile RAG systems capable of addressing a wide range of real-world user needs.
Acknowledgement
This work was supported by National Natural Science Foundation of China No. 62502049.
@@ Appendix A Implementation Details
In our evaluation of baseline methods on HawkRAG, we use BGE-M3 [Chen et al., 2023] as the retriever for vanilla RAG, RQ-RAG, HyDE, and MemoRAG, setting the hit number to 5. For methods that segment long contexts into chunks, we utilize the semantic-text-splitter tool, limiting chunks to a maximum of 512 tokens. MemoRAG employs the officially released memorag-qwen2-7b-inst as its memory model. For GraphRAG, we leverage GPT-4o for graph construction and use the retrieved context for generation. All baseline methods adopt Qwen-2.5-7B-instruct-128K as the generator.
@@
Appendix B Limitations
Appendix C Public Data and Model Memorization vs. Genuine Retrieval
Most publicly available web data, including domain-specific corpora, are likely included in the pre-training corpus of today’s large language models. This challenge is shared by nearly all modern NLP benchmarks.
Nevertheless, benchmarks built on such corpora remain meaningful for several reasons. First, seeing a text during pre-training does not guarantee full memorization, nor does it ensure accurate answers for queries requiring complex reasoning or synthesis. Our benchmark’s query–answer pairs are manually annotated to capture nuanced, multi-step information-seeking behaviors that go well beyond simple fact recall.
Second, to mitigate concerns regarding memorization versus genuine retrieval, we include evaluations using several strong commercial LLM APIs. By comparing the performance of a range of models on our benchmark, we can better assess the extent to which retrieval-augmented reasoning (rather than memorization) contributes to success. Specifically, we compare RAG methods with three leading commercial LLMs. The results are shown in Table 4.
The results demonstrate that while strong LLMs have memorized substantial information from public corpora, they still lag behind retrieval-augmented methods in overall performance. Notably, for factoid queries at Level 1 and Level 2, RAG methods outperform strong LLMs by a large margin, suggesting that even with exposure to the underlying texts during pre-training, LLMs cannot reliably recall fine-grained factual details. For Level 3 and Level 4 tasks, which require summarizing broad content or synthesizing information, strong LLMs perform comparably to RAG methods, as these queries demand less precise retrieval and more general reasoning.
In summary, these results show that even though portions of HawkBench may have been seen by strong LLMs during pretraining, it remains a robust benchmark for evaluating stratified RAG performance. Without retrieval, even advanced LLMs such as GPT-4.1 can only solve a small fraction of the tasks, highlighting the necessity of effective retrieval-augmented reasoning. Moreover, these experiments suggest that HawkBench not only provides a comprehensive testbed for RAG evaluation, but also serves as a tool for assessing the factual memorization capabilities of state-of-the-art LLMs.
Domain & Level gemini-2.5-flash gpt-4o-mini gpt-4.1 Best RAG (Qwen-2.5 7B)
Legal-Level 1 13.6 10.5 15.3 79.2
Legal-Level 2 25.2 20.6 30.0 45.9
Legal-Level 3 15.8 20.7 22.0 32.7
Legal-Level 4 12.4 13.2 14.3 17.9
Finance-Level 1 13.6 12.3 22.3 79.5
Finance-Level 2 25.2 13.3 23.1 54.7
Finance-Level 3 13.8 17.8 20.2 30.8
Finance-Level 4 13.4 15.8 18.9 20.4
Science-Level 1 13.6 12.2 13.2 45.1
Science-Level 2 25.1 17.8 21.2 33.8
Science-Level 3 15.7 21.4 20.1 27.3
Science-Level 4 13.4 17.2 17.4 20.4
Table 4: Comparison of commercial LLM APIs with the best-performing RAG system (Qwen-2.5 7B). RAG substantially outperforms LLMs in factoid queries (Level 1 and 2), while LLMs remain competitive in higher-level reasoning tasks (Level 3 and 4).
Appendix D Broader Impact
Our work aims to advance the robustness and generalizability of RAG systems by introducing a comprehensive benchmark, HawkBench, that stratifies tasks based on real-world information-seeking complexity. This can benefit a wide range of applications—such as question answering, legal and financial document analysis, and educational tutoring—by enabling more adaptive and reliable retrieval-augmented language models.
However, improving general-purpose information-seeking systems also raises concerns. These include the risk of propagating misinformation from retrieved content, amplifying biases present in training or retrieval corpora, and enabling misuse in sensitive domains without sufficient oversight. We encourage developers to adopt careful evaluation and safeguards when deploying RAG systems, especially in high-stakes or regulated scenarios.
Ultimately, we hope that HawkBench facilitates more transparent, equitable, and effective development of retrieval-based AI systems, while fostering research into more accountable and context-aware reasoning mechanisms.
@@ Appendix D Broader Impact
Table 5: Full details of main experimental results.